# Orbifold modifications of complex analytic varieties

### v4

@@ 4. Modification to double-point singularities
Let $X$ be a complex analytic variety. Let $S$ be a codimension 2 irreducible component of $X_{\operatorname{\textsubscript{\rm sing}}}$. Assume that there is a proper Zariski closed subset $C\subseteq S$ such that $X$ has the same type of canonical singularities at points of $S\setminus C$. Then there is a projective bimeromorphic morphism $f\colon Y\to X$ such that the following properties hold.
@@ 4.A. Hypersurface singularities
In this subsection, we consider the following situation. Let $X$ be a complex analytic variety and let $S\subseteq X$ be a closed subspace. Assume that $S$ is irreducible and smooth of dimension $n$. Let $W=\mathbb{D}^{N}\times S$. Assume that $X$ is isomorphic to the closed subspace in $W$ defined by holomorphic functions $f_{j}(x_{1},...,x_{N},s)$, where $(x_{1},...,x_{N})$ are the coordinates of $\mathbb{C}^{N}$, $s\in S$ is a point, and $j=1,...,k$ for some positive integer $k$. Assume that the composite inclusion $S\subseteq X\subseteq W$ coincides with $\{\mathbf{0}_{N}\}\times S\subseteq W$. In particular, $f_{j}(\mathbf{0}_{N},s)=0$ for all $s\in S$ and for all $j=1,...,k$. We set
@@
With the notation above, the embedding dimension of $X$ at a point $s\in S\subseteq X$ is equal to $N+n-r(s)$.
@@
Since $S$ is smooth, by Jacobi criterion (see for example [GR84, Section 6.1]), the embedding dimension of $X$ is equal to $N+n-R(s)$, where $R(s)$ is the rank of $(\mathrm{d}f_{j}(\mathbf{0}_{N},s))_{j=1,...,k}$, in the vector space $\Omega^{1}_{W,(\mathbf{0}_{N},s)}$. Since $f_{j}(\mathbf{0}_{N},s)=0$ for all $s\in S$ and for all $j=1,...,k$, the partial derivatives of $f_{j}$, with respect to the variables of $S$, are all zero along $\{\mathbf{0}_{N}\}\times S$, for all $j=1,...,k$. It follows that $R(s)$ is equal to $r(s)$. This completes the proof of the lemma. ∎
@@
With the notation above, let $r=\operatorname{min}\{r(s)\ |\ s\in S\}$. In other words, $N+n-r$ is the maximum of the embedding dimensions of $X$ at the points of $S$. Assume that $X$ has embedding dimension smaller than $N+n-r$ at general points of $S$. Then there is a projective bimeromorphic morphism $f\colon Y\to X$ obtained by blowing up $X$ at ideal sheaves whose cosupports are strictly contained in $S$, such that the maximum of the embedding dimensions of $Y$ at the points of $S_{Y}$ is at most $N+n-r-1$, where $S_{Y}$ is the strict transform of $S$ in $Y$. Moreover, $S_{Y}$ is smooth and the construction of $f$ is functorial, which depends only on the isomorphism class of $X$ around $S$. In particular, it is independent of $N$.
@@
Let $o\in S$ be a point at which the embedding dimension of $X$ is $N+n-r$. We will first work locally around $o$. In particular, we are allowed to shrink $S$ around $o$. Then, by showing that the operation in the lemma is functorial, we can glue the local constructions in the global setting.
@@
for $j=1,...,r.$ Then, by subtracting each $f_{j}$, with $j>r$, some linear combination, with coefficients as holomorphic functions on $S$, of the $f_{j}$’s with $j=1,...,r$, we may assume that
@@
for $i=1,...,r$, $j>r$ and all $s\in S$ contained in a neighborhood of $o$. In particular, the matrix $\Big{[}\frac{\partial f_{j}}{\partial x_{i}}(\mathbf{0}_{N},s)\Big{]}_{1\leq i\leq N,1\leq j\leq k}$, whose entries are holomorphic functions on $S$, is of the shape
@@
equivalently, the entries of the matrix $B$ above. Then, up to shrinking $S$, the closed subset contained in $S$, at which $X$ has embedding dimension $N+n-r$, is the cosupport of $\mathcal{I}$. By assumption, this cosupport is a proper subset of $S$.
Let $h\colon Z\to X$ be the projective bimeromorphic morphism induced by blowing up the ideal $\mathcal{I}$ on $S$, see Subsection 3.B. That is, $h$ is the blowup of $X$ at the ideal $(x_{1},...,x_{N},\mathcal{I})$. Let $S_{Z}$ be the strict transform of $S$ in $Z$ and $o_{Z}\in S_{Z}$ a point lying over $o$. Then there is a neighborhood $U$ of $o_{Z}$ in $Z$, which can be viewed as a closed subset of $\mathbb{D}^{N}\times(U\cap S_{Z}).$ Moreover, $h$ can be written in coordinates as
$h\colon(x_{1}^{\prime},...,x_{N}^{\prime},s^{\prime})\mapsto(tx_{1}^{\prime},...,tx_{N}^{\prime},h(s^{\prime})),$
where $t$ is a holomorphic function on $U\cap S_{Z}$ which defines the exceptional divisor of $S_{Z}\to S$ on $U\cap S_{Z}$. We define
@@
for $j=r+1,...,k$. Then as shown in Subsection 3.B, all of the $f^{\prime}_{j}$ are holomorphic functions and vanish on $Z$. We see that
@@
for $j=1,...,r.$ Furthermore, since the ideal $\mathcal{I}$ is generated by $\frac{\partial f_{j}}{\partial x_{i}}(\mathbf{0}_{N},s)$, where $i,j>r$, there are some $i,j>r$ such that
@@
As a consequence, after the blowup, the rank of $(\mathrm{d}_{x^{\prime}}f^{\prime}_{j}(\mathbf{0}_{N},o_{Z}))_{j=1,...,k}$ is at least $r(o)+1$. In particular, if $S_{Z}$ is smooth at $o_{Z}$, then the embedding dimension of $Z$ at $o_{Z}$ is at most $N+n-r-1$.
Let $f\colon Y\to X$ be the projective bimeromorphic morphism induced by the functorial principalization of $(g|_{S_{Z}})^{-1}\mathcal{I}\cdot\mathcal{O}_{S_{Z}}$ on $S_{Z}$. In particular, the strict transform $S_{Y}$ of $S$ in $Y$ is smooth. Since $Y\to Z$ is the composition of blowups with centers strictly contained in $S_{Z}$, the similar computation as above shows that $Y$ has embedding dimension at most $N+n-r-1$ at a point $o_{Y}\in S_{Y}$ lying over $o$.
@@
Therefore, if $(g_{e})_{e=1,...,t}$ is another collection of functions defining $X$, then they define the same ideal $\mathcal{I}$. Moreover, we assume that $\psi\colon(x^{\prime}_{1},...,x^{\prime}_{N},s)\to(x_{1},...,x_{N},s)$ is a change of coordinates of $\mathbb{C}^{N}\times S$. Then the $N\times N$ matrix
@@
Next we will reduce to the case when $r=0$, that is, the case when $N+n$ is the embedding dimension of $X$ at $o$. Indeed, from the previous discussions, thanks to the implicit function theorem, up to a change of coordinate of $\mathbb{C}^{N}\times S$, we may assume that $(f_{1},...,f_{r})=(x_{1},...,x_{r})$. For any $j=r+1,...,k$, we set
@@
$\varphi\colon(x_{1},...,x_{N},s)\mapsto(x_{1}^{\prime},...,x_{N}^{\prime},s^{\prime}),$
@@
Then when $a\neq 0$ or $b\neq 0$, and $x_{1}=x_{2}=x_{3}=x_{4}=0$, the rank of the Jacobi matrix of $(f_{1},f_{2},f_{3})$ is 2. At $\mathbf{0}_{6}$, the rank of the Jacobi matrix is 1. Then $X$ has embedding dimension 5 at this point. Now we blow up $\mathbb{C}^{6}$ at the ideal $(x_{1},x_{2},x_{3},x_{4},a,b)$ and obtain $h\colon W\to\mathbb{C}^{6}$. Then on the open set $U=W\setminus h^{-1}_{*}(\{a=0\})$, $h$ can be written in coordinate as
$h\colon(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime},x^{\prime}_{4},\alpha,\beta)\mapsto(\alpha x_{1}^{\prime},\alpha x_{2}^{\prime},\alpha x_{3}^{\prime},\alpha x^{\prime}_{4},\alpha,\alpha\beta).$
@@
With the notation in Lemma 4.4, we assume that $X$ has hypersurface singularities (respectively is smooth) around general points of $S$. Then there is a projective bimeromorphic morphism $f\colon X^{\prime}\to X$ such that the following properties hold.
@@
$X_{e}\to\cdots\to X_{1}=X,$
such that each $X_{i}\to X_{i+1}$ is the functorial operation in Lemma 4.4, and that the maximum of embedding dimensions of $X_{e}$ at points of $S_{e}$ is $\operatorname{dim}X+1$, where $S_{e}$ is the strict transform of $S$ in $X_{e}$. We let $X^{\prime}=X_{e}$ and let $f\colon X^{\prime}\to X$ be the composition of the sequence. Then it satisfies all the properties of the corollary. ∎
@@ 4.B. Double-point singularities
Let $X$ be a complex analytic variety of dimension $n$ and let $S\subseteq X_{\operatorname{\textsubscript{\rm sing}}}$ be an irreducible component of codimension $2$. Assume that $S$ is smooth, that $X$ has hypersurface singularities around $S$, and that $X$ has double-point singularities around general points of $S$. Then there is a projective bimeromorphic morphism $f\colon X^{\prime}\to X$ such that the following properties hold.
@@
Let $h\colon X^{\prime\prime}\to X$ be the blowup up of $X$ at the ideal $(x,y,z,\mathcal{I})$. As shown in Subsection 3.B, at every point $o^{\prime\prime}$ in the strict transform $S^{\prime\prime}$ of $S$ in $X^{\prime\prime}$, lying over $o$, there is an open neighborhood $U^{\prime\prime}\subseteq X^{\prime\prime}$ of $o^{\prime\prime}$ such that $U^{\prime\prime}$ is isomorphic to the hypersurface in $\mathbb{D}^{3}\times(S^{\prime\prime}\cap U^{\prime\prime})$ defined by the holomorphic function
@@
Let $X^{\prime}\to X^{\prime\prime}$ be the projective bimeromorphic morphism induced by the functorial principalization of the ideal $(h|_{S^{\prime\prime}})^{-1}\mathcal{I}\cdot\mathcal{O}_{S^{\prime\prime}}$ on $S^{\prime\prime}$, see Subsection 3.B. In particular, the strict transform $S^{\prime}$ of $S$ in $X^{\prime}$ is smooth. Since $X^{\prime}\to X^{\prime\prime}$ is the composition of blowups with center strictly contained in $S^{\prime\prime}$, the computations of Subsection 3.B implies that $X^{\prime}$ has double-point singularities along $S^{\prime}$.
To prove that construction is functorial, it is enough to show that the ideal $\mathcal{I}\subseteq\mathcal{O}(S)$ is independent of the choice of coordinates. Let $(\overline{x},\overline{y},\overline{z},\overline{s})$ be some other choice of coordinates of $\mathbb{D}^{3}\times S$ such that $X$ is defined by $\overline{F}(\overline{x},\overline{y},\overline{z},\overline{s})=0$. Let $\varphi\colon(x,y,z,s)\mapsto(\overline{x},\overline{y},\overline{z},\overline{s})$ be the change of coordinates, and we regard $\overline{x},\overline{y},\overline{z},\overline{s}$ as vector-valued holomorphic functions in $(x,y,z,s)$. Then we have
@@
Let $f\colon Y\to X$ be the normalization of $X$, and let $x\in X$ be a point. If $\mathcal{O}_{X,x}$ is an integral domain, then $f$ is a homeomorphism over an open neighborhood of $x$ by [GR84, Corollary on page 163]. Conversely, if $\mathcal{O}_{X,x}$ is not an integral domain, then there is an open neighborhood of $x$ which is not irreducible by the local decomposition lemma (see [GR84, page 79]). In particular, $f$ is not a homeomorphism over $x$.
Let $s\in S$ be a point. We claim that $\mathcal{O}_{X,s}$ is an integral domain. Assume the contrary. Then there is an open neighborhood $Z$ of $s$ in $X$ such that $Z$ is not irreducible. We may assume that $Z\subseteq\mathbb{D}^{n+1}$, and that $Z$ is a double cover of $\mathbb{D}^{n}$, where $n$ is the dimension of $X$. Then $Z$ has two irreducible components $Z_{1}$, $Z_{2}$, and each of them is a smooth divisor in $\mathbb{D}^{n+1}$. It follows that $Z_{\operatorname{\textsubscript{\rm sing}}}=Z_{1}\cap Z_{2}$, which is pure of dimension $n-1$. However, an irreducible component of $S\cap Z$ is an irreducible component of $Z_{\operatorname{\textsubscript{\rm sing}}}$, and is of dimension at most $n-2$. This is a contradiction.

### v6

@@ 4. Modification to double-point singularities
Let $X$ be a complex analytic variety. Let $S$ be a codimension 2 irreducible component of $X_{\operatorname{\textsubscript{\rm sing}}}$. Assume that there is a proper Zariski closed subset $C\subseteq S$ such that $X$ has the same type of canonical singularities at points of $S\setminus C$. Then there is a projective bimeromorphic morphism $f\colon Y\rightarrow X$ such that the following properties hold.
@@ 4.A. Hypersurface singularities
In this subsection, we consider the following situation. Let $X$ be a complex analytic variety and let $S\subseteq X$ be a closed subspace. Assume that $S$ is irreducible and smooth of dimension $n$. Let $W=\mathbb{D}^{N}\times S$. Assume that $X$ is isomorphic to the closed subspace in $W$ defined by holomorphic functions $f_{j}(x_{1},...,x_{N},s)$, where $(x_{1},...,x_{N})$ are the coordinates of $\mathbb{C}^{N}$, $s\in S$ is a point, and j=1,,kj=1,...,k for some positive integer $k$. Assume that the composite inclusion $S\subseteq X\subseteq W$ coincides with $\{\mathbf{0}_{N}\}\times S\subseteq W$. In particular, $f_{j}(\mathbf{0}_{N},s)=0$ for all $s\in S$ and for all j=1,,kj=1,...,k. We set
@@
With the notation above, the embedding dimension of $X$ at a point $s\in S\subseteq X$ is equal to N+nr(s)N+n-r(s).
@@
Since $S$ is smooth, by Jacobi criterion (see for example [GR84, Section 6.1]), the embedding dimension of $X$ is equal to N+nR(s)N+n-R(s), where $R(s)$ is the rank of $(\mathrm{d}f_{j}(\mathbf{0}_{N},s))_{j=1,...,k}$, in the vector space $\Omega^{1}_{W,(\mathbf{0}_{N},s)}$. Since $f_{j}(\mathbf{0}_{N},s)=0$ for all $s\in S$ and for all j=1,,kj=1,...,k, the partial derivatives of $f_{j}$, with respect to the variables of $S$, are all zero along $\{\mathbf{0}_{N}\}\times S$, for all j=1,,kj=1,...,k. It follows that $R(s)$ is equal to $r(s)$. This completes the proof of the lemma. ∎
@@
With the notation above, let $r=\operatorname{min}\{r(s)\ |\ s\in S\}$. In other words, N+nrN+n-r is the maximum of the embedding dimensions of $X$ at the points of $S$. Assume that $X$ has embedding dimension smaller than N+nrN+n-r at general points of $S$. Then there is a projective bimeromorphic morphism $f\colon Y\rightarrow X$ obtained by blowing up $X$ at ideal sheaves whose cosupports are strictly contained in $S$, such that the maximum of the embedding dimensions of $Y$ at the points of $S_{Y}$ is at most N+nr1N+n-r-1, where $S_{Y}$ is the strict transform of $S$ in $Y$. Moreover, $S_{Y}$ is smooth and the construction of $f$ is functorial, which depends only on the isomorphism class of $X$ around $S$. In particular, it is independent of $N$.
@@
Let $o\in S$ be a point at which the embedding dimension of $X$ is N+nrN+n-r. We will first work locally around $o$. In particular, we are allowed to shrink $S$ around $o$. Then, by showing that the operation in the lemma is functorial, we can glue the local constructions in the global setting.
@@
for j=1,,r.j=1,...,r. Then, by subtracting each $f_{j}$, with $j>r$, some linear combination, with coefficients as holomorphic functions on $S$, of the $f_{j}$’s with j=1,,rj=1,...,r, we may assume that
@@
for i=1,,ri=1,...,r, $j>r$ and all $s\in S$ contained in a neighborhood of $o$. In particular, the matrix $\Big[\frac{\partial f_{j}}{\partial x_{i}}(\mathbf{0}_{N},s)\Big]_{1\leq i\leq N,1\leq j\leq k}$, whose entries are holomorphic functions on $S$, is of the shape
@@
equivalently, the entries of the matrix $B$ above. Then, up to shrinking $S$, the closed subset contained in $S$, at which $X$ has embedding dimension N+nrN+n-r, is the cosupport of $\mathcal{I}$. By assumption, this cosupport is a proper subset of $S$.
Let $h\colon Z\rightarrow X$ be the projective bimeromorphic morphism induced by blowing up the ideal $\mathcal{I}$ on $S$, see Subsection 3.B. That is, $h$ is the blowup of $X$ at the ideal $(x_{1},...,x_{N},\mathcal{I})$. Let $S_{Z}$ be the strict transform of $S$ in $Z$ and $o_{Z}\in S_{Z}$ a point lying over $o$. Then there is a neighborhood $U$ of $o_{Z}$ in $Z$, which can be viewed as a closed subset of $\mathbb{D}^{N}\times(U\cap S_{Z}).$ Moreover, $h$ can be written in coordinates as
$h\colon(x_{1}^{\prime},...,x_{N}^{\prime},s^{\prime})\mapstochar\rightarrow(tx_{1}^{\prime},...,tx_{N}^{\prime},h(s^{\prime})),$
where $t$ is a holomorphic function on $U\cap S_{Z}$ which defines the exceptional divisor of $S_{Z}\rightarrow S$ on $U\cap S_{Z}$. We define
@@
for j=r+1,,kj=r+1,...,k. Then as shown in Subsection 3.B, all of the $f^{\prime}_{j}$ are holomorphic functions and vanish on $Z$. We see that
@@
for j=1,,r.j=1,...,r. Furthermore, since the ideal $\mathcal{I}$ is generated by $\frac{\partial f_{j}}{\partial x_{i}}(\mathbf{0}_{N},s)$, where $i,j>r$, there are some $i,j>r$ such that
@@
As a consequence, after the blowup, the rank of $(\mathrm{d}_{x^{\prime}}f^{\prime}_{j}(\mathbf{0}_{N},o_{Z}))_{j=1,...,k}$ is at least $r(o)+1$. In particular, if $S_{Z}$ is smooth at $o_{Z}$, then the embedding dimension of $Z$ at $o_{Z}$ is at most N+nr1N+n-r-1.
Let $f\colon Y\rightarrow X$ be the projective bimeromorphic morphism induced by the functorial principalization of $(g|_{S_{Z}})^{-1}\mathcal{I}\cdot\mathcal{O}_{S_{Z}}$ on $S_{Z}$. In particular, the strict transform $S_{Y}$ of $S$ in $Y$ is smooth. Since $Y\rightarrow Z$ is the composition of blowups with centers strictly contained in $S_{Z}$, the similar computation as above shows that $Y$ has embedding dimension at most N+nr1N+n-r-1 at a point $o_{Y}\in S_{Y}$ lying over $o$.
@@
Therefore, if $(g_{e})_{e=1,...,t}$ is another collection of functions defining $X$, then they define the same ideal $\mathcal{I}$. Moreover, we assume that $\psi\colon(x^{\prime}_{1},...,x^{\prime}_{N},s)\rightarrow(x_{1},...,x_{N},s)$ is a change of coordinates of $\mathbb{C}^{N}\times S$. Then the $N\times N$ matrix
@@
Next we will reduce to the case when $r=0$, that is, the case when $N+n$ is the embedding dimension of $X$ at $o$. Indeed, from the previous discussions, thanks to the implicit function theorem, up to a change of coordinate of $\mathbb{C}^{N}\times S$, we may assume that $(f_{1},...,f_{r})=(x_{1},...,x_{r})$. For any j=r+1,,kj=r+1,...,k, we set
@@
$\varphi\colon(x_{1},...,x_{N},s)\mapstochar\rightarrow(x_{1}^{\prime},...,x_{N}^{\prime},s^{\prime}),$
@@
Then when $a\neq 0$ or $b\neq 0$, and $x_{1}=x_{2}=x_{3}=x_{4}=0$, the rank of the Jacobi matrix of $(f_{1},f_{2},f_{3})$ is 2. At $\mathbf{0}_{6}$, the rank of the Jacobi matrix is 1. Then $X$ has embedding dimension 5 at this point. Now we blow up $\mathbb{C}^{6}$ at the ideal $(x_{1},x_{2},x_{3},x_{4},a,b)$ and obtain $h\colon W\rightarrow\mathbb{C}^{6}$. Then on the open set $U=W\setminus h^{-1}_{*}(\{a=0\})$, $h$ can be written in coordinate as
$h\colon(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime},x^{\prime}_{4},\alpha,\beta)\mapstochar\rightarrow(\alpha x_{1}^{\prime},\alpha x_{2}^{\prime},\alpha x_{3}^{\prime},\alpha x^{\prime}_{4},\alpha,\alpha\beta).$
@@
With the notation in Lemma 4.4, we assume that $X$ has hypersurface singularities (respectively is smooth) around general points of $S$. Then there is a projective bimeromorphic morphism $f\colon X^{\prime}\rightarrow X$ such that the following properties hold.
@@
$X_{e}\rightarrow\cdots\rightarrow X_{1}=X,$
such that each $X_{i}\rightarrow X_{i+1}$ is the functorial operation in Lemma 4.4, and that the maximum of embedding dimensions of $X_{e}$ at points of $S_{e}$ is $\operatorname{dim}X+1$, where $S_{e}$ is the strict transform of $S$ in $X_{e}$. We let $X^{\prime}=X_{e}$ and let $f\colon X^{\prime}\rightarrow X$ be the composition of the sequence. Then it satisfies all the properties of the corollary. ∎
@@ 4.B. Double-point singularities
Let $X$ be a complex analytic variety of dimension $n$ and let $S\subseteq X_{\operatorname{\textsubscript{\rm sing}}}$ be an irreducible component of codimension $2$. Assume that $S$ is smooth, that $X$ has hypersurface singularities around $S$, and that $X$ has double-point singularities around general points of $S$. Then there is a projective bimeromorphic morphism $f\colon X^{\prime}\rightarrow X$ such that the following properties hold.
@@
Let $h\colon X^{\prime\prime}\rightarrow X$ be the blowup up of $X$ at the ideal $(x,y,z,\mathcal{I})$. As shown in Subsection 3.B, at every point $o^{\prime\prime}$ in the strict transform $S^{\prime\prime}$ of $S$ in $X^{\prime\prime}$, lying over $o$, there is an open neighborhood $U^{\prime\prime}\subseteq X^{\prime\prime}$ of $o^{\prime\prime}$ such that $U^{\prime\prime}$ is isomorphic to the hypersurface in $\mathbb{D}^{3}\times(S^{\prime\prime}\cap U^{\prime\prime})$ defined by the holomorphic function
@@
Let $X^{\prime}\rightarrow X^{\prime\prime}$ be the projective bimeromorphic morphism induced by the functorial principalization of the ideal $(h|_{S^{\prime\prime}})^{-1}\mathcal{I}\cdot\mathcal{O}_{S^{\prime\prime}}$ on $S^{\prime\prime}$, see Subsection 3.B. In particular, the strict transform $S^{\prime}$ of $S$ in $X^{\prime}$ is smooth. Since $X^{\prime}\rightarrow X^{\prime\prime}$ is the composition of blowups with center strictly contained in $S^{\prime\prime}$, the computations of Subsection 3.B implies that $X^{\prime}$ has double-point singularities along $S^{\prime}$.
To prove that construction is functorial, it is enough to show that the ideal $\mathcal{I}\subseteq\mathcal{O}(S)$ is independent of the choice of coordinates. Let $(\overline{x},\overline{y},\overline{z},\overline{s})$ be some other choice of coordinates of $\mathbb{D}^{3}\times S$ such that $X$ is defined by $\overline{F}(\overline{x},\overline{y},\overline{z},\overline{s})=0$. Let $\varphi\colon(x,y,z,s)\mapstochar\rightarrow(\overline{x},\overline{y},\overline{z},\overline{s})$ be the change of coordinates, and we regard $\overline{x},\overline{y},\overline{z},\overline{s}$ as vector-valued holomorphic functions in $(x,y,z,s)$. Then we have
@@
Let $f\colon Y\rightarrow X$ be the normalization of $X$, and let $x\in X$ be a point. If $\mathcal{O}_{X,x}$ is an integral domain, then $f$ is a homeomorphism over an open neighborhood of $x$ by [GR84, Corollary on page 163]. Conversely, if $\mathcal{O}_{X,x}$ is not an integral domain, then there is an open neighborhood of $x$ which is not irreducible by the local decomposition lemma (see [GR84, page 79]). In particular, $f$ is not a homeomorphism over $x$.
Let $s\in S$ be a point. We claim that $\mathcal{O}_{X,s}$ is an integral domain. Assume the contrary. Then there is an open neighborhood $Z$ of $s$ in $X$ such that $Z$ is not irreducible. We may assume that $Z\subseteq\mathbb{D}^{n+1}$, and that $Z$ is a double cover of $\mathbb{D}^{n}$, where $n$ is the dimension of $X$. Then $Z$ has two irreducible components $Z_{1}$, $Z_{2}$, and each of them is a smooth divisor in $\mathbb{D}^{n+1}$. It follows that $Z_{\operatorname{\textsubscript{\rm sing}}}=Z_{1}\cap Z_{2}$, which is pure of dimension n1n-1. However, an irreducible component of $S\cap Z$ is an irreducible component of $Z_{\operatorname{\textsubscript{\rm sing}}}$, and is of dimension at most n2n-2. This is a contradiction.